Find the slope of the line whose equation is 7x - 2y = 4.
Found 2 solutions by malglu, Cromlix:
Answer by malglu(63)

rearrange to get
2y=7x-4
divide by 2
y=7/2x -2
the slope is 7/2

Answer by Cromlix(4381)

Sort out equation into y = mx + c
7x - 2y = 4
- 2y = -7x + 4
Multiply throughout by -1
2y = 7x - 4
y = 7/2x - 2
Slope of line = 7/2 or 3.5
Hope this helps.
:-)